# Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect

### v1

@@ Abstract
Introduction.– The Fulling-Davies-Unruh effect [1, 2, 3, 4] plays a crucial role in understanding the quantum fields in curved spacetime. A uniformly accelerated observer sees a thermal bath in the vacuum of an inertial observer. It also closely relates to the so-called Hawking radiation [5, 6] through the equivalence principle. The accelerated observer is analogous to the static observer far away from the black hole, and the static observer is analogous to the free-falling observer near the black hole horizon. In this way, detecting the Unruh effect backs the existence of the Hawking effect. Hence it may shed light on the quantum nature of strong gravity. However those phenomenon are far from experimental verification. Despite its importance, requirement of significantly high acceleration ($\sim 10^{21}m/s^{2}$ for $1$ Kelvin temperature [4]) hindering its experimental detection. Numerous efforts have been made to detect the Unruh effect in several approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However either the proposed linear acceleration is too high ($\sim 10^{9}m/s^{2}$ with qubit energy gap $\sim$ MHz [13]) or investigation has been done with circular motion (see e.g. [20] where required acceleration is reported as $\sim 10^{7}m/s^{2}$ with qubit energy gap $\sim 10$ MHz, rotating frequency $\sim 0.1$ MHz and radius of path $\sim 10^{-3}m$).
The adiabatic-cyclic evolution may generate a geometric phase (known as Pancharatnam-Berry phase (PBP)) in a quantum system [21, 22]. It contains information of systems’ motion and background environments [23, 24, 25, 26]. Also, interferometric set-ups require significantly less acceleration to detect the Unruh phenomenon using the phase [23, 24]. In this set-up, one employs the two Unruh-DeWitt (UD) detectors: one detector accelerates in one arm of the interferometer, and another inertially moves in the other arm. Both of the detectors accumulate some PBP depending on their motions. Detection of this phase difference can reveal an indirect observation of the Unruh effect. In free space, the phase difference is experimentally realisable with minimum acceleration $\sim 10^{17}m/s^{2}$ [23, 24].
Motivated by these facts here we aim to study PBP difference (denote as $\delta\phi_{B}$) between an linearly accelerated UD detector and the inertial one in presence of mirrors. The acceleration is along one of the directions parallel to mirror’s plane. We investigate two different situations: the detectors are in presence of (i) single mirror and (ii) in between double parallel mirrors, respectively. The first situation shows that when the detectors are at the same distance from the mirror, $\delta\phi_{B}$ is at the measurable range ($\gtrsim 5.27\times 10^{-6}\,rad$ [30]) for much lower accelerations depending on the distance ($z_{0}$) from the mirror and the atom’s energy gap. For the analysis, one may utilize tuneable flux qubits as detectors. With the energy gap of the qubit in GHz order and $z_{0}\sim 10^{1}-10^{4}\,m$, we observe a measurable phase difference with acceleration as low as $a\sim 10^{10}m/s^{2}$. For the double mirrors, we do similar set-ups. The atom is allowed to move within the space in between the mirrors. Here $\delta\phi_{B}$ depends on $z_{0}$, separation distance between the two mirrors $L$ and the atom’s energy gap. We can achieve measurable phase differences for acceleration as low as $a\sim 10^{9}m/s^{2}$ with suitable choices of parameters (e.g. $z_{0}\sim 10^{1}\,m$, $L\sim 10^{2}\,m$ and energy gap $\sim 1$ GHz). However qubits with energy gap in MHz order (as considered in [24, 13, 20]), we find the required acceleration for two mirrors brings down by considerable amount $a\sim 10^{6}m/s^{2}$. This is not only much less than the earlier proposals [24, 13, 20], but also is very much achievable within the current technology (e.g. using thermal gradient technique [31]).
Let us now proceed towards the calculation to justify our claim.
@@
The UD detector model and PBP.– Consider the total Hamiltonian of the system $\hat{H}=\hat{H}_{d}+\hat{H}_{\phi}+\hat{H}_{\text{int}}$. Here $\hat{H}_{d}$ is the Hamiltonian of the detector, is taken to be $\hat{H}_{d}=\frac{\hbar}{2}\omega_{0}\hat{\sigma}_{3}$, where $\hat{\sigma}_{3}$ is Pauli spin matrix and $\hbar\omega_{0}$ is the energy gap of the detector. $\hat{H}_{\phi}$ is the background field’s (real scalar field $\hat{\phi}$) Hamiltonian and $\hat{H}_{\text{int}}$ is the interaction Hamiltonian, given by $\hat{H}_{\text{int}}=\lambda\,\hat{m}\,\hat{\phi}$. Here, $\hat{m}$ is the operator (chosen to be as $\hat{m}=(1/2)\sum_{i}\hat{\sigma}_{i}$) representing the detector, which is analogous to the dipole operator corresponding to electro-magnetic interactions and $\lambda$ is the coupling constant. Initially, the field is considered to be in the Minkowski vacuum state, and the detector’s state is $\hat{\rho}(0)=|\psi(0)\rangle\langle\psi(0)|$. With perturbative analysis, the reduced density matrix at a later time can be found by solving the Kossakowski-Lindblad equation [32, 33, 34]
@@
where $a_{ij}=A(\delta_{ij}-\delta_{i3}\delta_{j3})-iB\epsilon_{ijk}\delta_{k3}$, and $\tau$ is the detector’s proper time.
@@
$\gamma(\omega_{0})=\frac{\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}}{\hbar^{2}}\int_{-\infty}^{\infty}d(\Delta\tau)\,e^{i\omega_{0}\Delta\tau}\,G_{W}(\Delta\tau)\,,$ (4)
@@
By construction, this green function has symmetry between two points $(z_{0},~{}L-z_{0})$. Note that consideration of only $n=0$ term of the above leads to the same for the single mirror situation, located at $z=0$. So in future whenever discussion on the single mirror case will be done, we will take only the $n=0$ term from the expressions, evaluated for double mirrors.
For our case, we find
@@
where $J(u)=\frac{\sin\left(\frac{2\omega_{0}c}{a}\sinh^{-1}(|\frac{au}{c^{2}}|)\right)}{|\frac{au}{c^{2}}|\sqrt{\left(\frac{au}{c^{2}}\right)^{2}+1}}$ and $\kappa=\frac{\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}}{\hbar c^{3}}$. Considering only $n=0$ term, one gets $A$ and $B$ in presence of single mirror. Discarding the $z_{0}$-dependent term from these latter $A$ and $B$, we obtain those for free space. For the double-mirror system the summation over $n$ goes from $\infty$- ∞ to $\infty$. However for very large values of $n$, (when $a(Ln+z_{0})\approx aLn$), two terms inside the summation will cancel each other ($J(Ln)-J(Ln+z_{0})\approx 0$). Therefore for sufficiently large values of $n$, we may set a cut off on upper and lower limit of $n$ in order to evaluate the summation in (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect). Hence in numerical calculation we choose max$|n|$ to a large finite value. In this set-up the possibility to measure minimum PBP ($\sim 10^{-6}~{}rad$) with low acceleration ($a$) can be realized as follows. For very low value of $a$ we have $\coth\frac{\pi\omega_{0}c}{a}\approx 1$ and hence (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect) implies $A\approx B$. Therefore in this situation the relations for $A$ and $B$ with emission rate $\gamma(\omega_{0})$ and absorption rate $\gamma(-\omega_{0})$ of the atom implies $\gamma(-\omega_{0})\approx 0$. Then $\delta\phi_{B}$ (see (5)) mostly depends on $\gamma(\omega_{0})$. Moreover $A(\approx B)$ is mostly dominated by the term contributed by the effects of mirror(s) (see the term with summation in (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect)). Therefore controlling suitable values of the parameters, like $z_{0}$, $L$, it may be possible to reach at the desired goal.
@@
Parameters towards detecting PBP.– We need to define dimensionless parameters for numerical analysis of the geometric phase. Denote dimensionless acceleration and distances as $a/\omega_{0}c$, $L\omega_{0}/c$ and $z_{0}\omega_{0}/c$, respectively. Therefore, one can see the quantities $\coth\left(\frac{\pi\omega_{0}c}{a}\right)$ and $J(Ln)$ (or $J(Ln+z_{0})$) are dimensionless. The constant $\kappa$ is also dimensionless, since $\lambda\hat{m}\hat{\phi}$ has to be dimensionally same as $\hbar\omega$ ($\hat{\phi}$ has dimension of $\sqrt{\hbar/\omega L^{3}}$). Thus, as expected phase $\phi_{B}$ is dimensionless.
@@
Our analysis aims to evaluate the effect of acceleration in spaces with or without mirrors. Therefore the object of our interest is $\delta\phi_{B}=\phi_{B}-\lim_{a\to 0}\phi_{B}$ for double, single and no-mirror systems. Also note that, in SI unit $\hat{H}_{d}\sim\hbar\omega_{0}\sim 10^{-25}$ with $\omega_{0}\sim 1$ GHz. In perturbative calculations, one must have $\hat{H}_{int}<\hat{H}_{d}$, or $\lambda\hat{m}\hat{\phi}<10^{-25}$. Alternatively, we can also write $\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}\langle\hat{\phi}\hat{\phi}\rangle<10^{-50}$. We know $G_{W}\sim\hbar/c\sim 10^{-42}$, thus $\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}<10^{-8}$ and hence $\kappa<10^{2}$. Therefore for the numerical analysis we choose $\kappa\approx 1$.
@@
In a currently achievable experimental set-up, one can measure a phase difference with precision as small as $5.27\times 10^{-6}\,rad$ [30]. Also, one can implement tunable flux qubits, which can have energy gap in the range 111110101010 GHz [37, 38, 39, 38, 40]. For our purpose, we use $\theta=\pi/4$ and $\omega_{0}=1$ GHz. With this choice, the dimensionless acceleration $a/\omega_{0}c=1$ implies $a=3\times 10^{17}m/s^{2}$, and $\omega_{0}z_{0}/c=1$ implies the distance $z_{0}=0.3m$. In Fig. 1, we have shown $\delta\phi_{B}$ double, single and no-mirror systems in solid, dashed and dotted lines, respectively. Free space $\delta\phi_{B}$ decays sharply around $a/\omega_{0}c=1$ (i.e., $a=3\times 10^{17}m/s^{2}$). At $a/\omega_{0}c=0.1$, $\delta\phi_{B}$ is in order $\sim 10^{-16}\,rad$; exceptionally below the experimentally observable range. However, in that acceleration, $\delta\phi_{B}$ for the double and single mirror systems are in the observable range. This indicates the possibility of observing the phase difference in much lower accelerations.
@@
Figure 2: We plotted $\delta\phi_{B}$ for single-mirror system with respect to $\omega_{0}z_{0}/c$ for (a) $a/\omega_{0}c=10^{-2}$, (b) $a/\omega_{0}c=10^{-3}$ and (c) $a/\omega_{0}c=10^{-7}$ (range of $\omega_{0}z_{0}/c$ is chosen where $\delta\phi_{B}$ has maximum values), respectively with the choice $\theta=\pi/4$.
[…]
Figure 3: We plotted $\delta\phi_{B}$ for double-mirror system with respect to $\omega_{0}z_{0}/c$ for (a) $L\omega_{0}/c=10.0$ and (b) $L\omega_{0}/c=500.0$ and $1000.0$, respectively with $\theta=\pi/4$.
@@
We next explore the same in presence of a double mirrors. In Fig. 3, we have plotted $\delta\phi_{B}$ with respect to $\omega_{0}z_{0}/c$ for $L\omega_{0}/c=10.0$ ($L=3m$), $500.0$ ($L=150m$) and $1000.0$ ($L=300m$), respectively. We choose $\omega_{0}=1$ GHz and $\theta=\pi/4$. The first subfigure with $L\omega_{0}/c=10.0$ shows that $\delta\phi_{B}$ is in the experimentally observable range when accelerations $a/\omega_{0}c\sim 10^{-5},~{}10^{-6}$ and $10^{-7}$ (i.e. $a\sim 10^{12},~{}10^{11}$ and $10^{10}m/s^{2}$, respectively). Second subfigure with $L\omega_{0}/c=500.0$ and $1000.0$ shows the similar nature with $a/\omega_{0}c\sim 10^{-8}$ ($a\sim 10^{9}m/s^{2}$). Both subfigures show that the highest peak of $\delta\phi_{B}$ decreases as we use lower values of accelerations. Note that for tiny accelerations, the value $\max|n|$ of the summation has to be much larger to satisfy $J(Ln)\approx J(Ln+z_{0})$. To find out the appropriate $\max|n|$ value, we checked the values of $\delta\phi_{B}$ for $\max|n|$ from $0.1$ to $1\,million$ (in intervals of $0.1\,million$) and used the particular $\max|n|$ value, after which $\delta\phi_{B}$ approximately remain constant. For our parameter values $a/\omega_{0}c\sim 10^{-5},~{}10^{-6},~{}10^{-7}$ and $10^{-8}$, we used value of $\max|n|=0.1,~{}0.4,~{}0.4$ and $0.5\,million$, respectively. In all the above cases, we used the energy gap of the qubit to be $1$ GHz. However for $\omega_{0}$ in MHz order (as mentioned in [24, 13, 20]), the required acceleration can be brought down up to $10^{6}m/s^{2}$ with $\omega_{0}=1$ MHz, $L=150$ (or $300\,km$), $z_{0}=3.3\,km$. This required acceleration is significantly lower than other configured systems in the literature [13].

### v2

# Mirror-enhanced acceleration-induced geometric phase

@@ Abstract
Fulling-Davies-Unruh effect contains great amount of theoretical importance in various branches of physics. Requirement of very high acceleration hinders its experimental evidence. We put forward an idea to experimentally probe this effect by utilizing the Pancharatnam-Berry phase of an accelerated atom in presence of mirrors. We show that for much lower accelerations, the phase gets significantly enhanced in the presence of mirrors. We propose a schematic design of an interferometric set-up to experimentally capture this effect by utilizing the phase difference between an accelerated and an inertial atoms. For the choice of hydrogen atoms and suitable separation between atoms and mirrors, the required acceleration can be very low.
Introduction.– The Fulling-Davies-Unruh effect [1, 2, 3, 4] plays a crucial role in understanding the quantum fields in curved spacetime. A uniformly accelerated observer sees a thermal bath in the vacuum of an inertial observer. Equivalence principle bridges it to Hawking radiation [5, 6]. The accelerated observer is analogous to the static observer far away from the black hole, and the static observer is analogous to the free-falling one near the black hole horizon. So Unruh effect detection backs the existence of Hawking effect. Hence it may shed light on the quantum nature of strong gravity. However those phenomena are far from experimental verification due to requirement of significantly high acceleration ($\sim 10^{21}m/s^{2}$ for $1$ Kelvin temperature [4]). Numerous efforts have been made to detect the effect [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However either the proposed linear acceleration is still too high ($\sim 10^{9}m/s^{2}$ with qubit energy gap $\sim$ MHz [13]) or investigation has been done with circular motion (see e.g. [20] where required acceleration is reported as $\sim 10^{7}m/s^{2}$ with qubit energy gap $\sim 10$ MHz, rotating frequency $\sim 0.1$ MHz and radius of path $\sim 10^{-3}m$).
The adiabatic-cyclic evolution may generate a geometric phase (known as Pancharatnam-Berry phase (PBP)) in a quantum system [21, 22]. It contains information of systems’ motion and background environments [23, 24, 25]. Also, interferometric set-ups require significantly less acceleration to detect the Unruh phenomenon using the phase [23]. In this set-up, one employs the two Unruh-DeWitt (UD) detectors: one detector accelerates in one arm of the interferometer, and another inertially moves in the other arm. Both of the detectors accumulate some PBP depending on their motions. Detection of this phase difference can reveal an indirect observation of the Unruh effect. In free space, the phase difference is experimentally realisable with minimum acceleration $\sim 10^{17}m/s^{2}$ [23, 26].
@@
Looking at the estimated parameters we propose a schematic experimental set-up. Choose the hydrogen atom (with energy gap $\sim$ GHz for $2p\to 2s$ [31]) as atoms to acquire measurable PBP ($\gtrsim 5.27\times 10^{-6}\,rad$ [30]). The atom can be accelerated by using various available techniques, e.g. temperature gradient technique [32], within two mirrors (e.g. the grounded conducting planes for electric field) which are separated by a distance $L\simeq 150~{}m$ along with $z_{0}\simeq 5~{}m$. The required acceleration $\sim 10^{9}~{}m/s^{2}$ is achieved by varying temperature of $1$ Kelvin within a distance $\sim 10\mu m$ using temperature gradient technique [32]. This proposed setup may give an idea to constructed the same at the lab.
@@
Model set-up. – Consider the total Hamiltonian of the system $\hat{H}=\hat{H}_{d}+\hat{H}_{\phi}+\hat{H}_{\text{int}}$. $\hat{H}_{d}$ is the Hamiltonian of the detector, is taken to be $\hat{H}_{d}=\frac{\hbar}{2}\omega_{0}\hat{\sigma}_{3}$, where $\hat{\sigma}_{3}$ is Pauli spin matrix and $\hbar\omega_{0}$ is the energy gap of the detector. $\hat{H}_{\phi}$ is the background field’s (real scalar field $\hat{\phi}$) Hamiltonian and $\hat{H}_{\text{int}}$ is the simplified light-matter interaction Hamiltonian, given by $\hat{H}_{\text{int}}=\lambda\,\hat{m}\,\hat{\phi}$. Here, $\hat{m}$ is the operator (chosen to be as $\hat{m}=\hat{\sigma}_{2}$) representing the detector, which is analogous to the dipole operator corresponding to electro-magnetic interactions and $\lambda$ is the coupling constant. Initially, the field is considered to be in the Minkowski vacuum state, and the detector’s state is $\hat{\rho}(0)=|\psi(0)\rangle\langle\psi(0)|$. With perturbative analysis, the reduced density matrix at a later time can be found by solving the Kossakowski-Lindblad equation [33, 34, 35]
@@
where $a_{ij}=A(\delta_{ij}-\delta_{i3}\delta_{j3})-iB\epsilon_{ijk}\delta_{k3}$, and $\tau$ is the detector’s proper time. The above one is valid under three approximations, namely Born, Markov and rotating-wave approximations [36].
@@
$\gamma(\omega_{0})=\frac{\lambda^{2}}{\hbar^{2}}\int_{-\infty}^{\infty}d(\Delta\tau)\,e^{i\omega_{0}\Delta\tau}\,G_{W}(\Delta\tau)\,,$ (4)
@@
Consideration of only $n=0$ term of the above leads to the same for the single mirror situation, located at $z=0$. So in future whenever discussion on the single mirror case will be done, we will take only the $n=0$ term from the expressions for double mirrors.
Here
@@
where $J(u)=\frac{\sin\left(\frac{2\omega_{0}c}{a}\sinh^{-1}(|\frac{au}{c^{2}}|)\right)}{|\frac{au}{c^{2}}|\sqrt{\left(\frac{au}{c^{2}}\right)^{2}+1}}$ and $\kappa=\frac{\lambda^{2}}{\hbar c^{3}}$. Considering only $n=0$ term, one gets $A$ and $B$ in presence of single mirror. Discarding the $z_{0}$-dependent term from these latter $A$ and $B$, we obtain those for free space. For the double-mirror system the summation over $n$ goes from $\infty$- ∞ to $\infty$. However for very large values of $n$, (when $a(Ln+z_{0})\approx aLn$), two terms inside the summation will cancel each other ($J(Ln)-J(Ln+z_{0})\approx 0$). Therefore for sufficiently large values of $n$, we may set a cut off on upper and lower limit of $n$ in order to evaluate the summation in (Mirror-enhanced acceleration-induced geometric phase). Hence in numerical calculation we choose max$|n|$ to a large finite value. The possibility to measure minimum PBP ($\sim 10^{-6}~{}rad$) with low acceleration ($a$) can be realized as follows. For very low value of $a$ we have $\coth\frac{\pi\omega_{0}c}{a}\approx 1$ and hence (Mirror-enhanced acceleration-induced geometric phase) implies $A\approx B$. Therefore in this situation the relations for $A$ and $B$ with emission rate $\gamma(\omega_{0})$ and absorption rate $\gamma(-\omega_{0})$ of the atom implies $\gamma(-\omega_{0})\approx 0$. Then $\delta\phi_{B}$ (see (5)) mostly depends on $\gamma(\omega_{0})$. Moreover $A(\approx B)$ is mostly dominated by the term contributed by the effects of mirror(s) (see the term with summation in (Mirror-enhanced acceleration-induced geometric phase)). Therefore controlling suitable values of the parameters, like $z_{0}$, $L$, it is possible to reach at the desired goal. The off-diagonal elements in (3) decay with time with decay constant $2A$. Moreover $A$ increases with the increase of $a$ which shows the on-set of decoherence.
@@
Searching feasible parameters.– We need to define dimensionless parameters for numerical analysis of the geometric phase. Denote dimensionless acceleration and distances as $a/\omega_{0}c$, $L\omega_{0}/c$ and $z_{0}\omega_{0}/c$, respectively. Therefore, one can see the quantities $\coth\left(\frac{\pi\omega_{0}c}{a}\right)$ and $J(Ln)$ (or $J(Ln+z_{0})$) are dimensionless. The constant $\kappa$ is also dimensionless, since $\lambda\hat{m}\hat{\phi}$ has to be dimensionally same as $\hbar\omega$ ($\hat{\phi}$ has dimension of $\sqrt{\hbar/\omega L^{3}}$). Thus, as expected phase $\phi_{B}$ is dimensionless.
@@
Our analysis aims to evaluate the effect of acceleration in spaces with or without mirrors. For this purpose, we use an interferometric set up, where one accelerated and one inertial atoms move inside the interferometer arms. The PB phase acquired by the accelerated atom is $\phi_{B}$, and the phase acquired by the inertial atom can be obtained as $\lim_{a\to 0}\phi_{B}$. The interference pattern will depend on the phase difference $\delta\phi_{B}=\phi_{B}-\lim_{a\to 0}\phi_{B}$, which is our the object of interest. Also note that, in SI unit $\hat{H}_{d}\sim\hbar\omega_{0}\sim 10^{-25}$ with $\omega_{0}\sim 1$ GHz. In perturbative calculations, one must have $\hat{H}_{int}<\hat{H}_{d}$, or $\lambda\hat{m}\hat{\phi}<10^{-25}$. Alternatively, we can also write $\lambda^{2}\langle\hat{\phi}\hat{\phi}\rangle<10^{-50}$. We know $G_{W}\sim\hbar/c\sim 10^{-42}$, thus $\lambda^{2}<10^{-8}$ and hence $\kappa<10^{2}$. Therefore for the numerical analysis we choose $\kappa\approx 1$. Following similar arguments, one will obtain $\kappa<1$ for $\omega_{0}\sim 100$ MHz. In this situation, we will still obtain detectable values of $\delta\phi_{B}$ (see [30]) for this choice of $\omega_{0}\sim 100$ MHz and $\kappa<1$, which we will explain later on.
@@
Currently the experimentally measurable phase difference is as small as $5.27\times 10^{-6}\,rad$ [30]. Also, hydrogen atoms ($2s\to 2p$ energy gap $\sim$ GHz [31]) can be considered as our atom system. Further choose $\theta=\pi/4$ and $\omega_{0}=1$ GHz. With this choice, $a/\omega_{0}c=1$ and $\omega_{0}z_{0}/c=1$ imply $a=3\times 10^{17}m/s^{2}$, and $z_{0}=0.3m$, respectively. Fig. 1, shows $\delta\phi_{B}$ for double, single and no-mirror systems in solid, dashed and dotted lines, respectively. In free space, $\delta\phi_{B}$ decays sharply around $\omega_{0}c/a=1$ (i.e., $a=3\times 10^{17}m/s^{2}$). At $\omega_{0}c/a=10$ (i.e., $a=3\times 10^{16}m/s^{2}$), $\delta\phi_{B}$ is in order $\sim 10^{-16}\,rad$; exceptionally below the experimentally observable range. However, in that acceleration, $\delta\phi_{B}$ for the double and single mirror systems are in the observable range. This indicates the possibility of observing the phase difference in much lower accelerations. We now explore the possibility.
@@
Figure 2: We plotted $\delta\phi_{B}$ for single-mirror system with respect to $\omega_{0}z_{0}/c$ for (a) $a/\omega_{0}c=10^{-3}$ and (b) $a/\omega_{0}c=10^{-7}$ (range of $\omega_{0}z_{0}/c$ is chosen where $\delta\phi_{B}$ has maximum values), respectively. $\delta\phi_{B}$ for double-mirror system with respect to $\omega_{0}z_{0}/c$ has been plotted for (c) $L\omega_{0}/c=10.0$ and (d) $L\omega_{0}/c=500.0$ and $1000.0$, respectively. Here we choose $\theta=\pi/4$.
In (a) and (b) of Fig. 2, we plot $\delta\phi_{B}$ Vs $\omega_{0}z_{0}/c$ for single mirror with $a/\omega_{0}c=10^{-3}$ and $10^{-7}$ (i.e. $\sim 10^{14}$ and $10^{10}m/s^{2}$), respectively. Observe that as we lower the acceleration, the highest peak of $\delta\phi_{B}$ appears for larger separation from the mirror ($z_{0}$). Also, the magnitude of $\delta\phi_{B}$ decreases as we use lower values of accelerations. We can lower the acceleration as long as $\delta\phi_{B}\gtrsim 5.27\times 10^{-6}\,rad$. Investigation shows the acceleration can be as low as $a/\omega_{0}c\sim 10^{-7}$ (or $a\sim 10^{10}m/s^{2}$) with $\omega_{0}z_{0}/c\approx 91000$ ($z_{0}=27.3km$) and the corresponding $\delta\phi_{B}\approx 1.1\times 10^{-5}\,rad$ (see subfigure (b) of Fig. 2).
@@
We next explore the same in presence of a double mirrors in (c) and (d) of Fig. 2. The plots have been done for $L\omega_{0}/c=10.0$ ($L=3m$), $500.0$ ($L=150m$) and $1000.0$ ($L=300m$), respectively. We choose $\omega_{0}=1$ GHz and $\theta=\pi/4$. The subfigure (c) with $L\omega_{0}/c=10.0$ shows that $\delta\phi_{B}$ is in the experimentally observable range when accelerations $a/\omega_{0}c\sim 10^{-5},~{}10^{-6}$ and $10^{-7}$ (i.e. $a\sim 10^{12},~{}10^{11}$ and $10^{10}m/s^{2}$, respectively). Subfigure (d) with $L\omega_{0}/c=500.0$ and $1000.0$ shows the similar nature with $a/\omega_{0}c\sim 10^{-8}$ ($a\sim 10^{9}m/s^{2}$). Both subfigures show that the highest peak of $\delta\phi_{B}$ decreases as we use lower values of accelerations. Note that for tiny accelerations, the value $\max|n|$ of the summation has to be much larger to satisfy $J(Ln)\approx J(Ln+z_{0})$. To find out the appropriate $\max|n|$ value, we checked the values of $\delta\phi_{B}$ for $\max|n|$ from $0.1$ to $1\,million$ (in intervals of $0.1\,million$) and used the particular $\max|n|$ value, after which $\delta\phi_{B}$ approximately remain constant (see Appendix). For our parameter values $a/\omega_{0}c\sim 10^{-5},~{}10^{-6},~{}10^{-7}$ and $10^{-8}$, we used value of $\max|n|=0.1,~{}0.4,~{}0.4$ and $0.5\,million$, respectively. In all the above cases, we used the energy gap of the atom to be $1$ GHz.
Since the value of the phase difference $\delta\phi_{B}$ is proportional to $\kappa$, reducing $\kappa$ will reduce the numerical value of $\delta\phi_{B}$. However for $\omega_{0}$ in 100 MHz order [14] with $\kappa<1$, $\delta\phi_{B}$ will remain in the experimentally observable range [30] where the required acceleration can be brought down up to $10^{8}m/s^{2}$ with $L=1.5$ (or $3\,km$), $z_{0}=33\,m$. This can be estimated following the earlier arguments. In this scenario, the numerical values of $\delta\phi_{B}$ will be rescaled and can be obtained from plot Fig. 2 where the values in $y$-axis now will correspond to ‘$\delta\phi_{B}/\kappa$’. On the other hand the corresponding values of ‘$\omega_{0}z_{0}/c$’ can be estimated from $x$-axis. Thus in this scenario, the required acceleration is significantly lower than other configured systems in the literature [13].
A schematic experimental set-up.– Searching for a possible experimental setup, as an example we consider hydrogen atom and concentrate on $2s\to 2p$ for which $\omega_{0}\sim$ GHz [31]. The atom is kept between two large mirrors (e.g., grounded conducting planes for electric field), separated at distance $L\approx 150m$ in the $z$-direction. An interferometric set-up is placed in between the mirrors, where two atoms move along the $x$-direction inside two arms of the interferometer before entering the interference slits. The atoms have an equal distance from the first mirror ($z_{0}\approx 5m$) and have a small separation in $y$-direction. For a proposal purpose, we accelerate the atom by using the thermal gradient technique ($k_{B}\Delta{T}=ma\Delta{x}$) [32, 39]. To reach an acceleration of $\sim 10^{9}m/s^{2}$ in a distance $\Delta{x}\sim 10\mu{m}$, the required temperature gradient (or change) is $\Delta{T}\sim 1$K. Here we consider $m\sim 10^{-27}kg$ ( typical mass of hydrogen atom). For practical purpose, the interaction time ($\Delta{t}$) is sufficient if $\Delta{t}>>\omega_{0}^{-1}(=10^{-9}s)$. Therefore, one sets $\Delta{t}\sim 1\mu{s}$, then final velocity is $v_{f}\sim 10^{3}m/s^{2}$, which comes under non-relativistic regime. The other atom moves with the same constant velocity $v_{f}$ but with a different time interval $\Delta{t}^{\prime}$ so that both atoms reach the slits simultaneously with identical velocities. To obtain this velocity, hydrogen (-like) atoms can be heated in a high-temperature oven. The atoms will have a distribution of velocities. However, by using velocity selectors (like rotating discs with slits which only allow atoms within a narrow velocity range to pass through), one can filter out atoms with the desired velocity [40, 41]. Moreover, the distance travelled by the atom is $\sim 1mm$. So length of the mirrors has to be much larger than this distance to neglect the edge effects of the mirrors. Also, the area of mirrors $\mathcal{A}$ needs to be such that $L^{2}<<\mathcal{A}$. With this proposed set-up, when the atoms pass through the slits one can detect the phase difference between the atoms form their interference pattern. This should be $\sim 10^{-5}rad$, which is much higher than the minimum experimentally measurable phase difference ($\gtrsim 5.27\times 10^{-6}\,rad$, see [30]). A schematic diagram of the experimental set-up is shown in Fig. 3. The bumps in the screen (interference pattern) are produced due to the phase difference between the beams (or, atoms). In usual interference experiments, one also observe the similar bumps due to path differences ($\Delta{\bar{x}}_{i}$). The constructive interference will occur if $\Delta{\bar{x}}_{i}=n\lambda$, where $\lambda$ is associated (de Broglie) wavelength of the atoms and $n$ is order of maxima point. Here $\lambda$ have to be equal for both particles, i.e., the velocity of the atoms near the slits have to be equal. $\Delta{\bar{x}}_{i}$ depends on the actual set-up and position of the maxima points on the screen. $\Delta{\bar{x}}_{i}$ is again related to the phase difference ($\Delta{\bar{\phi}}_{i}$) between the atoms as $\Delta{\bar{\phi}}_{i}=\frac{2\pi}{\lambda}\Delta{\bar{x}}_{i}$. Once we get the path difference from the interference pattern, we can calculate the associated phase difference. In this way, one can obtain the phase difference ($\Delta{\bar{\phi}}_{i}$) if both atoms are moving inertially (i.e., gained no additional phase difference due to the time evolution). However, in our specific set-up, we will have additional phase difference due to acceleration as PB phase difference. Therefore, the effective path difference ($\Delta{\bar{x}}_{eff}$) calculated from the interference pattern shall be different from $\Delta{\bar{x}}_{i}$. It shall also contain the contribution due to PB phase as $\Delta{\bar{x}}_{eff}=\Delta{\bar{x}}_{i}+\Delta{\bar{x}}_{PB}$, where $\Delta{\bar{x}}_{PB}$ is effective path difference associated to PB phase. $\Delta{\bar{x}}_{i}$ can be calculated from the usual set-up where both the atoms are inertial with same velocity. Whereas $\Delta{\bar{x}}_{eff}$ is obtained from the present proposed experimental setup. Then one can easily calculate $\Delta{\bar{x}}_{PB}\,(=\Delta{\bar{x}}_{eff}-\Delta{\bar{x}}_{i})$. Once we have $\Delta{\bar{x}}_{PB}$, we can obtain the experimental value of $\Delta{\bar{\phi}}_{PB}$ through $\Delta{\bar{\phi}}_{PB}=\frac{2\pi}{\lambda}\Delta{\bar{x}}_{PB}$. Then we shall compare this experimentally obtained phase with our theoretical predicted PB phase (also see [42, 43]).
Figure 3: A schematic diagram of the experimental set-up.